OpenStudy (anonymous):

What volume of 6.00 M HNO3 would be needed to completely react with 50.0g of Ca(OH)2 in the reaction below? Ca(OH)2(s) + 2HNO3(aq) ---> 2H2O(1) + Ca(NO3)2(aq)

OpenStudy (cuanchi):

1) calculate how many moles you have in 50.0g of Ca(OH)2 n=m/MM (number of moles equal to mass divided by the molecular mass) 2)By the stoichiometry of the reaction you can see that one mole of Ca(OH)2 will react with 2 moles of HNO3. Then you will need double amount of moles of HNO3 than moles of Ca(OH)2 calculated in the step 1. So, multiply the moles of Ca(OH)2 by 2 to find the moles of HNO3. 3) Calculate the volume of HNO3 M=n/V (molarity is equal to moles divided the volume in liters) V=n/M (divide the number of moles of HNO3 that you calculate in the step 2 by the concentration of the HNO3 given in the problem and you will have the volume in liters that you will need to react with the 50.0g of Ca(OH)2

OpenStudy (anonymous):

For step 1) I got 1.48 moles. For step 2) I multiplied by 2 to get 2.96 For step 3) I am confused

OpenStudy (cuanchi):

For step 1) I got 1.48 moles. NO!! 50.0/74.1=0.67

OpenStudy (cuanchi):

Ca=54.09 O=16 H=1